PART A Jabberwocky Inc. makes quarterly (end of period) payments of $10,000 into a pension fund earning 12 percent per y...

PART A

Jabberwocky Inc. makes quarterly (end of period) payments of $10,000 into a pension fund earning 12 percent per year compounded quarterly for 10 years. How much interest will they have earned in 10 years?

A. $40,000

B. 354,013

C. $550,255

D. $754,013

PART B

Joe Nautilus has $120,000 and wants to retire. What return must his money earn so he may receive annual benefits of $20,000 for the next 14 years.

A. Approx. 12%
B. Approx. 13%
C. Approx. 14%
D. Greater than 15%

PART C

Mr. Darden is selling his house for $165,000. He bought it for $75,000 nine years ago. What is the annual return on his investment?

A. 9.16%
B. 11.83%
C. 8.50%
D. None of the above
Answer #1

Part A

Option B $354013

B8 X fr =B7-B2 Α 1 Quarterly Payments 2 Total amount deposited 3 Annual Interest rate 4 Quarterly Interest Rate 5 Number of y

B8 - x fc =B7-B2 Α $ 10,000.00 $4,00,000.00 12% 3% 1 Quarterly Payments 2 Total amount deposited 3 Annual Interest rate 4 Qua

Part B

Option C Approx 14%

B4 - fc =RATE(B3,B2,-B1) B 1 Present Value 2 Annual Benefits 3 Time in years 4 Annual Rate to be earned 120000 20000 14 =RATE

B4 - : X fc =RATE(B3,B2,-B1) B 120000 20000 1 Present Value 2 Annual Benefits 3 Time in years 4 Annual Rate to be earned 14 1

Part C

Option A 9.16%

B4 - X ✓ fx =RATE(B3,0,-B2,B1) 165000 75000 1 Current Price 2 Buying Price 9 years ago 3 Time in years Rate Earned =RATE(B3,0

B4 X fx =RATE(B3,0,-B2,B1) B Α 1 Current Price 2 Buying Price 9 years ago 3 Time in years 4 Rate Earned 165000 75000 9.16%

The amount that is paid out is always entered with a negative sign while the received amount is with positive sign.